Consider the Christmas sales of the new BeCool Headsets e 12, which runs at $500 dollars and was released June. In 2018, surveyed owners of older BeCool products spent $1200 on their holiday shopping. (Five hundred dollars would be nearly half of their holiday budget for a single item.) In 2019, the 61 surveyed customers spent an average of $1220, with a standard deviation of $40. Has customer spending increased?   

**Fill page below from question**   

  EX 5: Hypothesis Testing: Read CANVAS assignment- THIRD PROBLEM: If GID ends odd, if GID ends even Kind of Variable: null val
Answer #1

5.

Kind of variable = Mean

null value = 1200

H0: \mu = 1200

H1: \mu > 1200

n = 61

\bar{x} = 1220

s = 40

Std. error = s/\sqrt{n}

Let \alpha = 0.05

Since the population mean is not known, we will use the t-test.

(3) Test Statistics The t-statistic is computed as follows: X - Mo t = 1220 – 1200 40/61. = 3.905 s/V%

T-Test Results: t-stats = 3.905, p-value =0.0001 0.00 -4.0 -3.5 -3.0 -2.5 -2.0 -1.5 -1.0 -0.5 0.0 0.5 1.0 2.5 3.0 3.5 14.0 1.

df = n-1

= 60

p-value = 0.0001

Since p-value = 0.0001 < 0.05 i.e. we can reject null hypohtesis and hence we can claim that the mean customer spending has  customer spending increased.

Please upvote if you have liked my answer, would be of great help. Thank you.